A monic quadratic is a quadratic in which the coefficient
of the quadratic term is 1. For example, r^2 - 3r + 7
is a monic quadratic, but 3t^2 - 3t + 1 is not.
A teacher writes a monic quadratic on the board.
Joanie copies the quadratic onto her paper, but writes
down the wrong constant term (but the correct quadratic
and linear terms).
She correctly factors the quadratic that she wrote down on
her paper, and determines that her quadratic has roots -16 and 2.
Kelvin copies the quadratic onto his paper, but he writes down
the wrong coefficient for the linear term (but the correct
quadratic and constant terms).
He correctly factors the quadratic that he wrote down, and
determines that his quadratic has roots -36 and 2.
What are the roots of the quadratic that the teacher wrote
on the board?

Answer by Edwin McCravy(20054)

A monic quadratic is a quadratic in which the coefficient
of the quadratic term is 1. For example, $r^2 - 3r + 7$
is a monic quadratic, but $3t^2 - 3t + 1$ is not.
A teacher writes a monic quadratic on the board.
The teacher wrote this:  x² + Ax + B

Joanie copies the quadratic onto her paper, but writes
down the wrong constant term (but the correct quadratic
and linear terms).
Joanie copied this:      x² + Ax + C

She correctly factors the quadratic that she wrote down on
her paper, and determines that her quadratic has roots -16 and 2.
So Joanie got this:                  x=-16; x=2
So Joanie had this before:          x+16=0; x-2=0 
So Joanie had this factorization:    (x+16)(x-2)
So Joanie copied:                    x²+14x-32  

Since Joanie had the right linear term, A = 14,
so the teacher wrote this:  x² + 14x + B 

Kelvin copies the quadratic onto his paper, but he writes down
the wrong coefficient for the linear term (but the correct
quadratic and constant terms).
Kevin copied this:   x² + Dx + B  

He correctly factors the quadratic that he wrote down, and
determines that his quadratic has roots -36 and 2.
So Kevin got this:                  x=-36; x=2
So Kevin had this before:           x+36=0, x-2=0 
So Kevin had this factorization:    (x+36)(x-2)
So Kevin copied:                    x²+34x-72  

Since Kevin had the right constant term, B = -72

So the teacher wrote this:  x² + 14x - 72 

What are the roots of the quadratic that the teacher wrote
on the board?
I'll leave that last part up to you!!!  Factor, set = 0, and 
solve for the values of x, which are the roots. 

Edwin